OpenStudy (anonymous):

graph the equation by plotting three points -3x+2y=2

OpenStudy (phi):

can you change it to y = mx+b form ? the first step is add+3x to both sides

OpenStudy (anonymous):

0x+5y=5

OpenStudy (phi):

no. just write +3x to both sides of -3x+2y=2

OpenStudy (anonymous):

-3x+3+2y+3=2+3

OpenStudy (phi):

closer.... write +3x (not +3) and when we add things to both sides, we do it only once per side (this is different from multiplying or dividing)

OpenStudy (anonymous):

+3x-3x+2y=2+3x

OpenStudy (phi):

yes. now notice on the left side 3x take away 3x is 0x or just 0 you get 2y= 2+3x normally we write the "x term" first.... we are adding and 2+3x is the same as 3x+2 (order can be switched) so 2y = 3x+2 now divide by 2 (when dividing, we have to divide *every* term)

OpenStudy (anonymous):

2y/2=3x/2+2/2 1y=1.5x+1

OpenStudy (phi):

yes. or just y = 1.5x +1 in geometry, we normally keep the slope 1.5 as a fraction. the slope 1.5 is the same as 1.5/1 change in y is 1.5 divided by change in x (1) but we don't want 1.5 because you can't plot decimals. so instead, write 1.5 as 3/2 so change in y (3 ) divided by change in x (2) are nice numbers. we have y = (3/2) x + 1 can you pick an x that gives a nice y ?

OpenStudy (anonymous):

1

OpenStudy (phi):

what is y when x is 1 ?

OpenStudy (anonymous):

do i put 1 in the equation

OpenStudy (phi):

yes...everywhere you see x, erase the x, and put in 1 instead and remember (3/2)x is short for (3/2) times x

OpenStudy (anonymous):

i got 2 1/2

OpenStudy (phi):

yes, so (1, 2 1/2) or (1,2.5) or (1, 5/2) (all are the same thing) but you don't like fractions or decimals. maybe a different x will give a nicer y ?

OpenStudy (anonymous):

2

OpenStudy (phi):

and x=2 means y = ?

OpenStudy (anonymous):

4

OpenStudy (phi):

so (2,4) is nice now use the slope 3/2 = change in y divided by change in x to find another point .

OpenStudy (anonymous):

1.5,0.5

OpenStudy (phi):

how did you get that ?

OpenStudy (anonymous):

3/2 and 2/4

OpenStudy (phi):

change in x means you add that to the x change in y means you add that to the y so if you are at (2,4) and the slope is 3/2 you know the change in x is the bottom number (the 2) and the top number is the change in y (the 3)

OpenStudy (anonymous):

so do i add 2+2 and 3+4

OpenStudy (phi):

yes

OpenStudy (anonymous):

4,7

OpenStudy (phi):

now you need one more point. you can also go backwards from (2,4) if you go -2 in the x direction, you go -3 in the y direction notice that (-3/-2 is the same slope as 3/2)

OpenStudy (anonymous):

yeah

OpenStudy (phi):

can you find a 3rd point ?

OpenStudy (anonymous):

wait what two points do i have so far

OpenStudy (phi):

2,4 and 4,7

OpenStudy (anonymous):

2,4 and 3,2

OpenStudy (anonymous):

ok

OpenStudy (phi):

definitely NOT 3,2 (if you put x=3 in the equation you will not get y=2)

OpenStudy (anonymous):

ok so to get my 3rd point do i do 2+4 and 3+7

OpenStudy (phi):

yes, that will work

OpenStudy (anonymous):

(2,4) (4,7) (6,10) are my three points?

OpenStudy (phi):

if you plot them and draw a line through them you should get a nice line

OpenStudy (anonymous):

ok thank you can you please help me with one more problem

OpenStudy (phi):

notice the x values go 2,4,6 up by 2 and the y values go 4,7,10 up by 3 that means the slope is 3/2 (change in y divided by change in x is 3/2 ) notice if you go backwards: x down by 2, y down by 3 you get (0,1)

OpenStudy (anonymous):

yeah

OpenStudy (phi):

post the next one.